OpenStudy (anonymous):

geometric where: t6=4/27 and t9=4/729 find the first three terms

OpenStudy (anonymous):

4/27*r^3 = 4/729 r =1/3 a*r^5 = 4/27 a= 36 b= 12 c =4

OpenStudy (anonymous):

but the answer is 36, 12, 4

OpenStudy (anonymous):

okay i get it now thanks